the mean weight of baggage of 25 passengers from usa bound for the philippines is 31.8 kilos with a variance of 16. test the claim that the population mean baggage weight is equal to 30 kilos at 0.01 level of significance. Assume that the population is normally distributed. The t value is 2.492




1
Expert's answer
2022-03-29T17:49:24-0400

"H_0:\mu =30\\H_1:\mu \ne 30\\T=\sqrt{n}\frac{\bar{x}-\mu}{s}=\sqrt{25}\frac{31.8-30}{\sqrt{16}}=2.25\\Since\,\,\left| T \right|<2.492, the\,\,null\,\,hypothesis\,\,is\,\,not\,\,declined. \mu =30."